# Finiteness of formal pushforwards

### v1

@@ 1. Introduction
Theorem (Theorem 6.6).
Let $\mathscr{X}$ be a normal connected quasi-projective $T$-scheme, and let $f:V\hookrightarrow U$ be an inclusion of non-empty open subsets of the reduced closed fiber of $\mathscr{X}$ such that the complement of $V$ in $U$ has codimension at least two in $U$. Write $\mathfrak{U},\mathfrak{V}$ for the formal completions along $U,V$. If $\mathscr{F}$ is a torsion-free coherent sheaf on $\mathfrak{V}$, then $\widehat{f}_{*}(\mathscr{F})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@
Structure of the manuscript: We provide background and context in Section 2, followed by two commutative algebra results in Section 3, and general results on formal schemes and formal patches in Section 4. Using that material, in Section 5 we obtain a key result (Proposition 5.3) that asserts that the intersection of two finitely generated torsion-free formal modules is also finitely generated under a codimension two hypothesis on the complement of the union. In Section 6 we first show that for formal schemes, as for schemes, pushforward preserves quasi-coherence. Afterwards we obtain Theorem 6.6, mentioned above, in which the key property to prove is finiteness. A version of that result with a stronger conclusion is proven in Section 7 in the situation in which the modules are assumed to be flat, rather than just being torsion-free; see Proposition 7.4. Finally, in Section 8, patching problems are discussed, and Corollaries 8.3 and 8.4 shown.
@@ 2. Background and context
If $f:V\to U$ is a quasi-compact and quasi-separated morphism of schemes (e.g., an inclusion of Noetherian schemes), and if $\mathcal{F}$ is quasi-coherent on $V$, then $f_{*}(\mathcal{F})$ is quasi-coherent on $U$ (see [Sta25, Lemma 01LC]). For coherent modules, there is the following result, which is known to the experts, and which is essentially a special case of [EGA4, Partie 2, Corollaire 5.11.4(ii)] and [Sta25, Lemma 0AWA] (as Johan de Jong pointed out to us). Note that this theorem holds in particular in the case mentioned in the introduction, viz., of a normal connected quasi-projective scheme $U$, since normal (and integral) schemes are reduced, and since quasi-projective varieties are excellent (by [Mat80, Section 34]).
@@ 6. Formal pushforwards
Let $T$ be a complete discrete valuation ring with uniformizer $t$, and let $\mathscr{X}$ be a normal integral $T$-scheme of finite type, with reduced closed fiber $X$. Let $g:V\to U$ be an inclusion of open subsets of $X$, with inclusion $\widehat{g}:\mathfrak{V}\to\mathfrak{U}$ of the associated formal schemes.
@@
Let $\mathscr{X},U_{0},U_{1},U_{2},U,W$ be as in Proposition 5.3, and let $\widehat{g}:\mathfrak{W}\to\mathfrak{U}$ be the inclusion of formal schemes associated to $W,U$. Let $\mathscr{N}$ be a torsion-free coherent sheaf on the formal scheme $\mathfrak{W}$ that is associated to $W$. Then $\widehat{g}_{*}(\mathscr{N})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@
Let $T$ be a complete discrete valuation ring with residue field $k$, and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$. Let $f:V\to U$ be an inclusion of open subsets of $X$ such that $V$ is dense in $U$, and write $\mathfrak{V},\mathfrak{U}$ for the formal schemes associated to $V,U$. Let $\mathscr{F}$ be a torsion-free coherent sheaf on $\mathfrak{U}$. Then $\mathscr{F}$ is a subsheaf of $\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})$ via the natural morphism $\mathscr{F}\to\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})$.
@@
Let $T$ be a complete discrete valuation ring and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$. Let $U$ be a non-empty open subset of $X$, and let $V$ be an open subset of $U$ whose complement in $U$ has codimension at least two. Let $f:V\to U$ be the inclusion map, and write $\mathfrak{U},\mathfrak{V}$ for the formal schemes associated to $U,V$. Let $\mathscr{F}$ be a torsion-free coherent sheaf on $\mathfrak{V}$. Then $\widehat{f}_{*}(\mathscr{F})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@ 7. The flat case
Proposition 7.4.
Let $T$ be a complete discrete valuation ring with residue field $k$, and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$. Let $f:V\hookrightarrow U$ be an inclusion of connected open subsets of $X$ such that the complement of $V$ in $U$ has codimension at least two, and write $\mathfrak{V},\mathfrak{U}$ for the formal schemes associated to $V,U$. Let $\mathscr{F}$ be a flat coherent sheaf on $\mathfrak{U}$. Then $\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})=\mathscr{F}$.
@@
Concerning the necessity of the connectivity hypothesis in Proposition 7.4, consider the situation in Remark 7.2(b), and let $V=U_{1}\cup U_{2}$ there, with inclusion morphism $f:V\to U$. Then the complement of $V$ in $U$ (viz., the point $P$) has codimension at least two in $U$, and $V$ is disconnected. Let $\mathfrak{U},\mathfrak{V}$ be the formal completions of $U,V$, with inclusion map $\widehat{f}:\mathfrak{V}\to\mathfrak{U}$. Then $\widehat{f}_{*}\widehat{f}^{*}(\mathcal{O}_{\mathfrak{U}})$ is strictly bigger than $\mathcal{O}_{\mathfrak{U}}$, since the former sheaf is “doubled” at the point $P$ (corresponding to the strict containment $\widehat{R}_{U}\subset\widehat{R}_{U_{1}}\cap\widehat{R}_{U_{2}}$ in Remark 7.2(b)).

### v2

@@ 1. Introduction
Theorem (see Theorem 6.6).
Let $\mathscr{X}$ be a normal connected quasi-projective $T$-scheme, and let $f:V\hookrightarrow U$ be an inclusion of non-empty open subsets of the reduced closed fiber of $\mathscr{X}$ such that the complement of $V$ in $U$ has codimension at least two in $U$. Write $\mathfrak{U},\mathfrak{V}$ for the formal completions of $\mathscr{X}$ along $U,V$. If $\mathscr{F}$ is a torsion-free coherent sheaf on $\mathfrak{V}$, then $\widehat{f}_{*}(\mathscr{F})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@
Structure of the manuscript: We provide background and context in Section 2, followed by two commutative algebra results in Section 3, and general results on formal schemes and formal patches in Section 4. Using that material, in Section 5 we obtain a key result (Proposition 5.3) that asserts that the intersection of two finitely generated torsion-free formal modules is also finitely generated under a codimension two hypothesis on the complement of the union. In Section 6 we first show that for formal schemes, as for schemes, pushforward preserves quasi-coherence. Afterwards we obtain Theorem 6.6, mentioned above, in which the key property to prove is finiteness. A version of that result with a stronger conclusion is proven in Section 7 in the situation in which the modules are assumed to be flat, rather than just being torsion-free; see Theorem 7.4. Finally, in Section 8, patching problems are discussed, and Corollaries 8.3 and 8.4 shown.
@@ 2. Background and context
If $f:V\to U$ is a quasi-compact and quasi-separated morphism of schemes (e.g., an inclusion of Noetherian schemes), and if $\mathcal{F}$ is quasi-coherent on $V$, then $f_{*}(\mathcal{F})$ is quasi-coherent on $U$ (see [Sta25, Lemma 01LC]). For coherent modules, there is the following result, which is known to the experts, and which is essentially a special case of [EGA4, Partie 2, Corollaire 5.11.4(ii)] and [Sta25, Lemma 0AWA] (as Johan de Jong pointed out to us). Note that this theorem holds in particu1lar in the case mentioned in the introduction, viz., of a normal connected quasi-projective scheme $U$, since normal (and integral) schemes are reduced, and since quasi-projective varieties are excellent (by [Mat80, Section 34]).
@@ 6. Formal pushforwards
Let $T$ be a complete discrete valuation ring with uniformizer $t$, and let $\mathscr{X}$ be a normal integral $T$-scheme of finite type, with reduced closed fiber $X$ and formal completion $\mathfrak{X}$. Let $g:V\to U$ be an inclusion of open subsets of $X$, with inclusion $\widehat{g}:\mathfrak{V}\to\mathfrak{U}$ of the associated formal open subschemes of $\mathfrak{X}$.
@@
Let $\mathscr{X},U_{0},U_{1},U_{2},U,W$ be as in Proposition 5.3, let $\mathfrak{X}$ be the formal completion of $\mathscr{X}$, and let $\widehat{g}:\mathfrak{W}\to\mathfrak{U}$ be the inclusion of the formal open subschemes of $\mathfrak{X}$ that are associated to $W,U$. Let $\mathscr{N}$ be a torsion-free coherent sheaf on the formal scheme $\mathfrak{W}$ that is associated to $W$. Then $\widehat{g}_{*}(\mathscr{N})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@
Let $T$ be a complete discrete valuation ring with residue field $k$, and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$ and formal completion $\mathfrak{X}$. Let $f:V\to U$ be an inclusion of open subsets of $X$ such that $V$ is dense in $U$, and write $\mathfrak{V},\mathfrak{U}$ for the formal open subschemes of $\mathfrak{X}$ associated to $V,U$. Let $\mathscr{F}$ be a torsion-free coherent sheaf on $\mathfrak{U}$. Then $\mathscr{F}$ is a subsheaf of $\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})$ via the natural morphism $\mathscr{F}\to\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})$.
@@
Let $T$ be a complete discrete valuation ring and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$ and formal completion $\mathfrak{X}$. Let $U$ be a non-empty open subset of $X$, and let $V$ be an open subset of $U$ whose complement in $U$ has codimension at least two. Let $f:V\to U$ be the inclusion map, and write $\mathfrak{U},\mathfrak{V}$ for the formal open subschemes of $\mathfrak{X}$ associated to $U,V$. Let $\mathscr{F}$ be a torsion-free coherent sheaf on $\mathfrak{V}$. Then $\widehat{f}_{*}(\mathscr{F})$ is a torsion-free coherent sheaf on $\mathfrak{U}$.
@@ 7. The flat case
Theorem 7.4.
Let $T$ be a complete discrete valuation ring with residue field $k$, and let $\mathscr{X}$ be a quasi-projective normal integral $T$-scheme with reduced closed fiber $X$ and formal completion $\mathfrak{X}$. Let $f:V\hookrightarrow U$ be an inclusion of connected open subsets of $X$ such that the complement of $V$ in $U$ has codimension at least two, and write $\mathfrak{V},\mathfrak{U}$ for the formal open subschemes of $\mathfrak{X}$ associated to $V,U$. Let $\mathscr{F}$ be a flat coherent sheaf on $\mathfrak{U}$. Then $\widehat{f}_{*}\widehat{f}^{*}(\mathscr{F})=\mathscr{F}$.
@@
Concerning the necessity of the connectivity hypothesis in Theorem 7.4, consider the situation in Remark 7.2(b), and let $V=U_{1}\cup U_{2}$ there, with inclusion morphism $f:V\to U$. Then the complement of $V$ in $U$ (viz., the point $P$) has codimension at least two in $U$, and $V$ is disconnected. Let $\mathfrak{U},\mathfrak{V}$ be the formal completions of $U,V$, with inclusion map $\widehat{f}:\mathfrak{V}\to\mathfrak{U}$. Then $\widehat{f}_{*}\widehat{f}^{*}(\mathcal{O}_{\mathfrak{U}})$ is strictly bigger than $\mathcal{O}_{\mathfrak{U}}$, since the former sheaf is “doubled” at the point $P$ (corresponding to the strict containment $\widehat{R}_{U}\subset\widehat{R}_{U_{1}}\cap\widehat{R}_{U_{2}}$ in Remark 7.2(b)).
Let $f:V\hookrightarrow U$ be an inclusion of connected open subsets of the reduced closed fiber $X\subset\mathscr{X}$ as in Theorem 7.4, where the complement of $V$ in $U$ has codimension at least two. If $\mathscr{G}$ is a flat coherent sheaf on the formal scheme $\mathfrak{V}$ associated to $V$, the pushforward $\widehat{f}_{*}(\mathscr{G})$ is a torsion-free coherent sheaf on $\mathfrak{U}$ by Theorem 6.6, and we can ask whether it is flat. But in fact, flatness for such formal pushforwards need not hold in dimension at least three, even in the regular case, and similarly in the context of schemes (rather than formal schemes). In particular, there is the following example.
Example 7.6.
  1. (a)
    We first consider the scheme case. In [OSS11, Example 1.1.13], the authors give an example of a coherent sheaf $F$ on $X=\mathbb{P}^{3}_{\mathbb{C}}$ that is reflexive (i.e., agrees with its double dual $F^{\vee\vee}$) but is not locally free (so not flat), though it is locally free away from a certain closed point $x_{0}$. Let $V$ be the complement of $x_{0}$ in $X$, and let $G=F|_{V}$. Thus $G$ is locally free. Write $f:V\to X$ for the inclusion map, so that $G=f^{*}F$. Since $F$ is reflexive, and since the complement of $V$ has codimension at least two (in fact, three), it follows from [Hts80, Proposition 1.6] that $F\cong f_{*}G$. Thus $G$ is a flat coherent sheaf on $V$, but $f_{*}G$ is not flat.
  2. (b)
    We use the above example to produce an example in the case of formal schemes. Preserving the above notation, take $\mathscr{X}=\mathbb{P}^{3}_{\mathbb{C}[[t]]}$, and consider the coherent sheaf $\mathscr{F}:=\pi^{*}(F)$ on $\mathscr{X}$, where $\pi:\mathscr{X}\to X=\mathbb{P}^{3}_{\mathbb{C}}$ is the morphism induced by the inclusion $\mathbb{C}\to\mathbb{C}[[t]]$. Observe first that $\mathscr{F}$ is a reflexive sheaf on $\mathscr{X}$; this follows from the fact that if $M,N$ are modules over a ring $R$ with $M$ finitely presented, and if $S$ is a flat $R$-algebra, then the natural map $S\otimes_{R}M^{\vee}\to(S\otimes_{R}M)^{\vee}$ is an isomorphism (see [Eis95, Proposition 2.10]). Note also that $\mathscr{F}$ is not locally free, since if it were then it would still be locally free (hence flat) modulo $(t)$. Moreover the restriction $\mathscr{G}$ of $\mathscr{F}$ to $\mathscr{V}:=V\times_{\mathbb{C}}\mathbb{C}[[t]]$ is locally free on $\mathscr{V}$, since the pullback of a free module is free. Let $\mathfrak{X}$ be the formal scheme associated to $\mathscr{X}$, let $\mathfrak{V}$ be the formal subscheme of $\mathfrak{X}$ associated to $V\subset X$, and let ^$\widehat{\mathscr{F}},\widehat{\mathscr{G}}$ be the induced formal coherent sheaves on $\mathfrak{X},\mathfrak{V}$. Then $\widehat{\mathscr{G}}=\widehat{f}^{*}(\widehat{\mathscr{F}})$ is flat because it is the completion of the finitely generated flat $\mathcal{O}_{\mathscr{X}}$-module $\mathscr{G}$; and $\widehat{\mathscr{F}}$ is not flat because it is not locally free (since its closed fiber is not locally free).
    By Lemma 6.5, $\widehat{\mathscr{F}}$ is a subsheaf of $\widehat{f}_{*}\widehat{f}^{*}(\widehat{\mathscr{F}})$. To show that $\widehat{\mathscr{G}}$ is an example of a flat formal coherent sheaf on $\mathfrak{V}$ whose pushforward $\widehat{f}_{*}(\widehat{\mathscr{G}})$ to $\mathfrak{X}$ is not flat, it remains to prove that ^$\widehat{\mathscr{F}}=\widehat{f}_{*}\widehat{f}^{*}(\widehat{\mathscr{F}})=\widehat{f}_{*}(\widehat{\mathscr{G}})$. To do this, it suffices to show that for every affine open subset $U$ of $X$, we have ^$\widehat{\mathscr{F}}(U)=\widehat{f}_{*}\widehat{f}^{*}\widehat{\mathscr{F}}(U)$. This is trivial if $x_{0}\not\in U$; so we assume $x_{0}\in U$, and write $U^{\prime}=U\cap V=f^{-1}(U)$, the complement of $x_{0}$ in $U$. The restriction map ^$\widehat{\mathscr{F}}(U)\to\widehat{\mathscr{F}}(U^{\prime})$ is given by the inclusion ^$\widehat{\mathscr{F}}(U)\subseteq\widehat{f}_{*}\widehat{f}^{*}\widehat{\mathscr{F}}(U)=\widehat{\mathscr{F}}(U^{\prime})$; and our goal is now to show that this is an isomorphism.
    To do this, first note that the inclusion $\mathcal{O}_{\mathfrak{X}}(\mathfrak{U})\hookrightarrow\mathcal{O}_{\mathfrak{X}}(\mathfrak{U}^{\prime})$ is an isomorphism, by applying Theorem 7.4 to the sheaf $\mathcal{O}_{\mathfrak{U}}=\mathcal{O}_{\mathfrak{X}}|_{U}$. Since $U$ is affine and $\widehat{\mathscr{F}}$ is a coherent formal sheaf, the restriction ^$\widehat{\mathscr{F}}|_{U}$ is of the form $M^{\Delta}$ for some finitely generated $\mathcal{O}_{\mathfrak{X}}(\mathfrak{U})$-module $M$ (by [EGA1, Propositions 10.10.5, 10.10.2(i)]); and so the natural map ^$\widehat{\mathscr{F}}(U)\to F(U)$ via reduction modulo $(t)$ is surjective. But $F(U)=F(U^{\prime})$, since $F\cong f_{*}G$ by part (a). Thus the reduction map ^$\widehat{\mathscr{F}}(U^{\prime})\to F(U^{\prime})$ modulo $(t)$ is surjective, using that ^$\widehat{\mathscr{F}}(U)\subseteq\widehat{f}_{*}\widehat{f}^{*}\widehat{\mathscr{F}}(U)=\widehat{\mathscr{F}}(U^{\prime})$. So ^$\widehat{\mathscr{F}}(U)+(t)\widehat{\mathscr{F}}(U^{\prime})=\widehat{\mathscr{F}}(U^{\prime})$. Also, the module ^$\widehat{\mathscr{F}}(U^{\prime})$ is finitely generated over $\mathcal{O}_{\mathfrak{X}}(\mathfrak{U}^{\prime})=\mathcal{O}_{\mathfrak{X}}(\mathfrak{U})$, and the ideal $(t)$ is contained in the radical of $\mathcal{O}_{\mathfrak{X}}(\mathfrak{U}^{\prime})$. So by a version of Nakayama’s Lemma (see [Mat80, Corollary to Lemma 1.M]), ^$\widehat{\mathscr{F}}(U)=\widehat{\mathscr{F}}(U^{\prime})=\widehat{f}_{*}\widehat{f}^{*}\widehat{\mathscr{F}}(U)$, as desired. So indeed ^$\widehat{\mathscr{F}}=\widehat{f}_{*}\widehat{f}^{*}\widehat{\mathscr{F}}$, and $\widehat{\mathscr{G}}$ provides the asserted example.
But as we now show, if we restrict attention to the dimension two case, then there is a flatness assertion for pushforwards, in both the scheme and formal scheme situations.
Proposition 7.7.
  1. (a)
    Let $X$ be an excellent (e.g., quasi-projective) regular scheme of dimension two, and let $V\subseteq X$ be an open subset whose complement has codimension two in $X$. Write $f:V\to X$ for the inclusion map. If $G$ is a flat coherent sheaf on $V$, then $f_{*}(G)$ is a flat coherent sheaf on $X$.
  2. (b)
    Let $T$ be a complete discrete valuation ring, and let $\mathscr{X}$ be a two-dimensional regular quasi-projective flat $T$-scheme with reduced closed fiber $X$ and formal completion $\mathfrak{X}$. Let $f:V\hookrightarrow U$ be an inclusion of open subsets of $X$ such that the complement of $V$ in $U$ has codimension two, and write $\mathfrak{V},\mathfrak{U}$ for the formal open subschemes of $\mathfrak{X}$ associated to $V,U$. Let $\mathscr{G}$ be a flat coherent sheaf on $\mathfrak{V}$. Then $\widehat{f}_{*}(\mathscr{G})$ is a flat coherent sheaf on $\mathfrak{U}$.
Proof.
In part (a), $F:=f_{*}(G)$ is a torsion-free coherent sheaf on $X$ by Theorem 2.1, since $G$ is a torsion-free coherent sheaf on $V$. Also, $F$ is flat over $V$ since $G$ is; and so it remains to check flatness at the (isolated) points of $X$ in the complement of $V$. Since flatness is local, we may assume that $X$ is the spectrum of a two-dimensional regular local ring, and that $V$ is the complement of the closed point $P$. Here $V\subset X$ is the only strict inclusion $U^{\prime}\subset U$ of open subsets of $X$ such that the complement of $U^{\prime}$ in $U$ has codimension two. Moreover $f_{*}(F|_{V})=f_{*}f^{*}F=f_{*}f^{*}f_{*}(G)=f_{*}(G)=F$, since $f^{*}f_{*}(G)=G$. So by [Hts80, Proposition 1.6], $F$ is reflexive. Since $X$ is regular of dimension two, it then follows from [Hts80, Corollary 1.4] that $F$ is flat.
For part (b), Theorem 6.6 says that $\mathscr{F}:=\widehat{f}_{*}(\mathscr{G})$ is a torsion-free coherent sheaf on $\mathfrak{U}$. We wish to show that it is flat. Take a closed point $P\in U$; we will show that $\mathscr{F}$ is free on an open neighborhood of $P$ in $\mathfrak{U}$. Let $G,F$ be the restrictions of $\mathscr{G},\mathscr{F}$ to the reduced closed fiber $X$. Thus $G$ is flat on $V$, and $F=f_{*}(G)$. By part (a), $F$ is flat, and hence free on an affine open neighborhood $W$ of $P$ in $U$, say of rank $n$. After shrinking $W$, we may assume that there is an affine open neighborhood $\mathscr{W}=\operatorname{Spec}(R)$ of $P$ in $\mathscr{X}$ that meets $X$ in $W$. Since $\mathcal{O}_{\mathscr{X},P}$ is a regular local ring, it is a UFD, and every height one prime is principal. So after shrinking $W$ again, we may assume that the closed subscheme $W\subset\mathscr{W}$ is defined by a principal ideal $I=(s)$ for some element $s\in R$; and this element is regular because $\mathscr{X}$ is flat over $T$. Let $\mathfrak{W}$ be the formal open subscheme of $\mathfrak{X}$ associated to $W$. Thus $\mathcal{O}_{\mathfrak{X}}(W)=\mathcal{O}_{\mathfrak{X}}(\mathfrak{W})$ is the $I$-adic completion $\widehat{R}$ of $R$. Also, $\mathcal{O}_{X}(W)=\bar{R}:=R/I=\widehat{R}/I\widehat{R}$ is reduced, since $X$ is.
By the freeness of $F$ on $W$, we may choose an isomorphism $\bar{R}^{n}\to F(W)=\mathscr{F}(\mathfrak{W})/I\mathscr{F}(\mathfrak{W})$. Let $\bar{a}_{1},\dots,\bar{a}_{n}$ be the images of the standard basis elements of $\bar{R}^{n}$, and choose lifts $a_{1},\dots,a_{n}$ of the elements $\bar{a}_{i}$ to $\mathscr{F}(\mathfrak{W})$. We thus obtain a lift of the above isomorphism to a homomorphism $\widehat{R}^{n}\to\mathscr{F}(\mathfrak{W})$, taking the standard basis elements of $\widehat{R}^{n}$ to the lifts $a_{i}$. We claim that this map is injective. To see this, take some non-zero element $(r_{1},\dots,r_{n})\in\widehat{R}^{n}$ in the kernel. Since $\mathscr{F}$ is torsion-free, we may divide the elements $r_{i}$ by any common factor that is a power of $s$, and still have an element in the kernel. So we may assume that some $r_{i}$ is not divisible by $s$. But then the image $(\bar{r}_{1},\dots,\bar{r}_{n})\in\bar{R}^{n}$ of $(r_{1},\dots,r_{n})$ is non-zero but is in the kernel of the isomorphism $\bar{R}^{n}\to F(W)$. This is a contradiction, proving the claim.
Thus the image of $\widehat{R}^{n}\to\mathscr{F}(\mathfrak{W})$ is a free $R$-submodule $N$ of $\mathscr{F}(\mathfrak{W})$ with the property that $N+I\mathscr{F}(\mathfrak{W})=\mathscr{F}(\mathfrak{W})$. Since $I\widehat{R}$ is contained in the radical of $\widehat{R}$ and $\mathscr{F}(\mathfrak{W})$ is a finitely generated $\widehat{R}$-module, $N=\mathscr{F}(\mathfrak{W})$ by [Mat80, Corollary to Lemma 1.M]. Thus $\mathscr{F}(\mathfrak{W})$ is free, as desired. ∎